OpenStudy (anonymous):

The sum of two numbers, a and b, is 12. The first number, a, is 8 more than the second number

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

HOW R U CONFUSED

OpenStudy (jim766):

first step is to translate it a + b = 12 sum of two numbers, a& b, is 12 a = b + 8 1st number a is 8 more than 2nd now I would use substitution to solve. Can you finish it?

OpenStudy (anonymous):

so it'd be like 12 = (b + 8) + b [since a = b+ 8] right? combine like terms: 12 = 2b + 8 use transposition 12 - 8 = 2b [therefore: 4 = 2b ] then divide by 2 [ 4/2 = 2b/2] so b = 2; now substitute b = 2 to find a [12 = a + 2] should I also answer that?

OpenStudy (calculusxy):

Shouldn't it be 4 because 4+8=12

OpenStudy (calculusxy):

It has to be 2+10 because 10 is eight greater than 2.

OpenStudy (calculusxy):

Or 10+2

OpenStudy (anonymous):

no, @calculusxy, it is not 4 since the other term is 8 and remember we stated before that 8 - 4 should be 8? well, it is not, so we will have to look to other numbers then! But yes it is 10 :) simply put, we explain this by getting the other term: 2 (since 10 + 2 is indeed 12) and we notice that it also fits the other requirement: 10 - 2 = 8 :) so yay! you solved it :) hope this helped! have a great day :)